OpenStudy (anonymous):

Amy has deposited $661 in a savings account that earns interest at a rate of 3.1% compounded monthly. What will the account balance be in 15 years?

OpenStudy (kropot72):

Account balance in 15 years will be \[661(1+\frac{0.031}{12})^{15\times 12}\]=\[661(1+\frac{0.031}{12})^{180}\]